Based on the frequency distribution above, find the relative frequency for the class 35-38


AgesNumber of students
15-186
19-2210
23-2610
27-302
31-346
35-384

Based on the frequency distribution above, find the relative frequency for the class 35-38. 

Relative Frequency = _______ %  

Give your answer as a percent, rounded to one  decimal place
Answer #1

Total number of students = 6+10+10+2+6+4 = 38

The frequency of students from ages 35-38 is 4.

Relative frequency for the class 35-38 =

= (Frequency of students from 35-38 / Total students)*100

= (4/38)*100

= 10.5% (rounded to one decimal place).

IF YOU FIND THIS ANSWER HELPFUL, PLEASE UP VOTE.